# The birth mass function of neutron stars

### v1

@@ Abstract
The birth mass function of neutron stars encodes rich information about supernova explosions, double star evolution, and properties of matter under extreme conditions. To date, it has remained poorly constrained by observations, however. Applying probabilistic corrections to account for mass accreted by recycled pulsars in binary systems to mass measurements of 90 neutron stars, we find that the birth masses of neutron stars can be described by a unimodal distribution that smoothly turns on at $\mathbf{1.1\,\mathrm{\mathrm{M}_{\odot}}}$, peaks at $\mathbf{\approx 1.27\,\mathrm{\mathrm{M}_{\odot}}}$, before declining as a steep power law. Such a “turn-on” power-law distribution is strongly favoured against the widely-adopted empirical double-Gaussian model at the $\mathbf{3\sigma}$ level. The power-law shape may be inherited from the initial mass function of massive stars[1, 2], but the relative dearth of massive neutron stars implies that single stars with initial masses greater than $\mathbf{\approx 18\,\mathrm{\mathrm{M}_{\odot}}}$ do not form neutron stars, in agreement with the absence of massive red supergiant progenitors to supernovae[3, 4].
@@
Henan Academy of Sciences, Zhengzhou 450046, Henan, China
@@
Department of Physics and Synergetic Innovation Center for Quantum Effects and Applications, Hunan Normal University, Changsha, Hunan 410081, China
@@
Institute of Interdisciplinary Studies, Hunan Normal University, Changsha, Hunan 410081, China
@@
Centre for Gravitational Astrophysics, College of Science, The Australian National University, ACT 2601, Australia
@@
Institute for Globally Distributed Open Research and Education (IGDORE)
@@
Australia Telescope National Facility, CSIRO Space and Astronomy, P.O. Box 76, Epping, NSW 1710, Australia
@@
The observational determination of the neutron star mass function is a long-standing problem in astrophysics. Early attempts were only able to derive loose constraints on the range of possible neutron star masses from a small number of measurements[5, 6]. For a long time, all observed neutron stars mases were in a narrow range, consistent with a Gaussian distribution with a mean of $1.35\,\mathrm{\mathrm{M}_{\odot}}$ and a width of $0.04\,\mathrm{\mathrm{M}_{\odot}}$[7]. Due partly to the precise mass measurements of the Hulse-Taylor binary pulsar[8, 9], routinely a canonical mass of $1.4\,\mathrm{\mathrm{M}_{\odot}}$ is adopted in most current studies of neutron stars and in textbooks.
@@
As the number of neutron star mass measurements steadily increased, especially with the discoveries of massive neutron stars around $2\,\mathrm{\mathrm{M}_{\odot}}$[10, 11], it became apparent that the single-Gaussian model could no longer fit all available data. Various groups started to employ a two-Gaussian model to explain the observations, with the secondary mass peak shifting to higher values as more massive neutron stars are measured[12, 13, 14, 15]. The most widely adopted model consists of a narrow peak at $1.4\,\mathrm{\mathrm{M}_{\odot}}$ and a broad peak at $1.8\,\mathrm{\mathrm{M}_{\odot}}$[16], and the two peaks are attributed to different formation channels and evolutionary history of neutron stars[17].
@@
From a theoretical perspective, the neutron star mass function is a powerful probe into supernova physics, compact binary evolution, and the neutron star equation of state. It was once proposed that two low-mass peaks (at $1.25\,\mathrm{\mathrm{M}_{\odot}}$ and $1.35\,\mathrm{\mathrm{M}_{\odot}}$) claimed in the mass distribution are the product of distinct supernova progenitor channels, namely, electron capture supernovae and low-mass iron core collapse supernovae[18, 19]. An additional high-mass peak centred around $1.8\,\mathrm{\mathrm{M}_{\odot}}$ in the mass distribution was thought to arise due to the transition from convective central carbon burning to radiative central carbon burning[18, 20]. Recent detailed studies, however, reveal a complicated landscape of interacting shells with discontinuous changes in the core and remnant masses when the number of shells changes, or shells merge, as a function of initial or helium core mass[21, 23].
@@
Efforts to link the observed mass distribution to neutron star formation channels are complicated by the fact that neutron stars in binary systems often go through a recycling process and gain some mass through accretion[24, 25, 26, 27]. Such a recycling process is the standard formation pathway for millisecond pulsars, which have by far dominated the population of neutron stars with precise mass measurements. Whereas the recycling process may increase the neutron star mass by up to $\sim 0.3$ M[24, 27, 28], it was argued that the high-mass peak around $1.8\,\mathrm{\mathrm{M}_{\odot}}$ in the observed mass spectrum is probably not caused by mass accretion, but instead a hint of neutron stars being born massive[15]. Current supernova simulations, however, have difficulties in producing enough neutron stars around $1.8\,\mathrm{\mathrm{M}_{\odot}}$ that implied in the empirical two-Gaussian model[29].
@@
To determine the neutron star mass distribution, we compile a sample of 90 neutron stars for which well-determined mass estimates are available from observations of radio pulsars, gravitational waves and X-ray binaries. For the majority of neutron stars in our sample, the observed spin properties allow us to securely categorize neutron stars into two subclasses: recycled and nonrecycled (slowly rotating) neutron stars. The first class is characterised by rapid spins (period $P\lesssim 100\,\mathrm{ms}$), while the second class includes neutron stars with spin periods on the order of seconds or longer. For the purpose of constraining the birth mass function of neutron stars, the key difference between the two subclasses is that observed masses of recycled pulsars need to be corrected for mass accreted throughout the recycling process, whereas the measured masses of slow neutron stars should equal their birth masses, since no mass-gaining process occurred. See Section A, Extended Data Tables 1 and 2 in Methods for details.
@@
Figure 1 shows the distributions of measured masses (black lines with grey shaded area) and estimated birth masses (blue lines) for 39 recycled pulsars. In each subplot, we also list the pulsar name, observed spin period and plausible initial spin period. The birth masses of recycled pulsars are estimated by subtracting accreted masses from their measured masses. We adopt an analytical approach to estimate the accreted mass; to the leading order, the total amount of accreted mass scales inversely with the final equilibrium spin period in the recycling process. Based on the simple accretion spin-up theory[24, 30, 27], we derive the probability distribution of the accreted mass by accounting for uncertainties in the accretion rate, neutron star moment of inertia, magnetic inclination angle, and accretion disk-magnetosphere interactions. See Section B in Methods.
@@
We fit a variety of models to the mass distribution of neutron stars. In addition to the Gaussian-family models, we consider a turn-on power law distribution, which allows an exponential turn-on from the minimum mass[31], as well as a skewed Student’s t-distribution. Some of the models are illustrated in the bottom panel of Figure 2. Bayesian model selection is performed in order to find the model that best describes the data. The Bayes factors, which quantify the statistical support for one model against another, between five top-performing models and the empirical two-Gaussian (2G) model are shown in the top panel of Figure 2. The best-performing model is the turn-on power law (TOP), which is preferred against the 2G model with a Bayes factor greater than 300. Assuming equal prior probability for both models, this result means an odds of $0.003$, or equivalently $\approx 3\sigma$ significance.
@@
Figure 3 shows the reconstructed neutron star birth mass function for the turn-on power law model, as well as histograms for two different bin sizes. We find that the minimum mass is $1.1^{+0.04}_{-0.05}\,\mathrm{\mathrm{M}_{\odot}}$, and the distribution peaks at $1.27^{+0.03}_{-0.04}\,\mathrm{\mathrm{M}_{\odot}}$ and declines with a power-law index of $6.5^{+1.3}_{-1.2}$. We find marginal evidence for a maximum-mass cutoff, with a Bayes factor of 2. The maximum mass is loosely constrained to be $2.36^{+0.29}_{-0.17}\,\mathrm{\mathrm{M}_{\odot}}$. (All numbers quoted are median and $1\sigma$ credible intervals.) The posterior distributions of model parameters are presented in Figure 4. We find no evidence for additional peaks on top of the power-law function, although such a possibility cannot be ruled out either.
@@
The final evolution of the lowest-mass stars that make iron core collapse supernovae is remarkably diverse, and is a challenge to model due to the complicated physics[32]. Hence, any constraints from observations on this mass regime are invaluable. Such low-mass iron core collapse supernovae are indeed predicted to produce neutron stars with masses as low as or potentially below $1.15\,\mathrm{M}_{\odot}$[18, 32], consistent with our observed minimum. Electron-capture supernovae are unlikely to produce such low-mass neutron stars[33]. They, however, may make neutron stars with masses within the peak of the distribution. The current data does not provide support for two discrete low-mass peaks that are expected from theoretical calculations[19]; instead, the distribution is likely a superposition of several overlapping discontinuous (as a function of initial mass) contributions[21, 23], overlaid with some scatter due to variations in the turbulent supernova explosions[34]. Particularly marked is the absence of the second predicted peak at $1.35\,\mathrm{\mathrm{M}_{\odot}}$, likely hinting at the larger importance of lower-mass core collapse and of stripped core supernova progenitors[35].
@@
If we accept the core structures predicted by current stellar evolution models, the absence of prominent peaks and the measured power-law slope of the neutron star mass distribution has important implications for supernova explosion physics. Specifically, the most parsimonious explanation of the turn-on power law is that neutron stars from more massive progenitors undergo substantial net accretion during the explosion, and that explosions only occur for stars with main sequence masses of $\lesssim 18\,\mathrm{\mathrm{M}_{\odot}}$ in agreement with the observed absence of massive red supergiants as progenitors for core-collapse supernovae[3, 4], or for stripped stars with similar core structures. We demonstrate this through the use of semi-analytic supernova models[21], as described in Methods Section F.
Knowledge of the neutron star mass function is also essential to understand the population of binary neutron star mergers being observed by their gravitational wave signal. It allows us to obtain a more accurate prediction of the frequency band for post-merger gravitational wave signals, and can significantly benefit the design and optimization of future-generation gravitational-wave observatories that aim to probe post-merger physics[37, 38, 39]. The remnant compact object formed in a binary neutron star merger is in an extreme, high-density, high-energy state and encodes unique information about nuclear matter that the progenitor stars and the inspiral phase do not provide. Observations of the post-merger gravitational waves will provide an unprecedented opportunity to explore hot equations of state at supranuclear densities and identify phase transitions in extreme nuclear matter[40, 41, 42]. The post-merger emissions are at frequencies mainly determined by the progenitor neutron star masses and the yet unknown nuclear equation of state, spanning a wide range roughly from $1\,\mathrm{kHz}$ to $4\,\mathrm{kHz}$[43, 44].
@@
Figure 1: Individual mass distributions of 39 recycled pulsars. Black lines with grey shaded area are measured masses, while blue lines are estimated birth masses after subtracting accreted masses based on the analytical prescription. Observed and plausible initial spin periods, along with pulsar names are given in each subplot. For readability, the X-label name in ${\rm M}_{\odot}$ is omitted and we scale different posteriors to the same Y-coordinate height.
@@
Figure 2: Top: the Bayes factor between various models (M) and the two-Gaussian (2G) model for the birth mass function of neutron stars. The models are: TOP – turn-on power law distribution; $\rm{2G^{max}_{min}}$ – two-Gaussian distribution with both a low- and high-mass cutoff; SST – skewed Student’s t distribution; $\rm{2G_{min}}$ – two-Gaussian distribution with a low-mass cutoff; $\rm{2G^{max}}$ – two-Gaussian distribution with a high-mass cutoff. Bottom: Illustrations of mass models. In addition to models listed in the top panel, we also show the single-Gaussian distribution with a low and high-mass cutoff ($\rm{G^{max}_{min}}$). Model parameters are set at their maximum-likelihood values inferred from data.
@@
Figure 3: The reconstructed birth mass function of neutron stars based on the preferred turn-on power-law (TOP) model. The solid blue line shows the population predictive distribution, while the shaded blue band indicates the 90% credible interval. Also plotted are a 10-bin histogram with error bars (90% credibility) and a 20-bin histogram for the data, showing the robustness of our model in fitting the underlying distribution.
@@
Figure 4: The posterior distribution of the turn-on power-law model parameters, including the minimum neutron star mass $m_{\rm min}$, the peak of the mass function $m_{\rm peak}$, the maximum mass $m_{\rm max}$, and the power-law index $\alpha$.
@@
We thank Zhenwei Li, Xuefei Chen and Zhanwen Han for providing data from their numerical simulations that are plotted in Extended Data Figure 2. We thank Ilya Mandel for useful discussions. This work was supported by the National Natural Science Foundation of China under grant numbers 12203004 (XZ), 12305059 (ZQY), 12405056 (ZCC), 12021003 (ZHZ). XZ and HG are supported by the Fundamental Research Funds for the Central Universities. This work was supported in part by the Australian Research Council (ARC) Centre of Excellence for Gravitational Wave Discovery (OzGrav), through project numbers CE170100004 and CE230100016.. SS is a recipient of an ARC Discovery Early Career Research Award (DE220100241). This work was supported by ARC Discovery grants DP240101786 (BM, AH) and DP240103174 (AH).
Supplementary Information is linked to the online version of the paper at http://www.nature.com/nature.
@@
Reprints and permissions information is available at http://www.nature.com/reprints. The authors declare no competing financial interests. Correspondence and requests for materials should be addressed to zhuxj@bnu.edu.cn and zhuzh@bnu.edu.cn.

### v2

@@ Abstract
The birth mass function of neutron stars encodes rich information about supernova explosions, double-star evolution and properties of matter under extreme conditions. To date, it has remained poorly constrained by observations, however. Applying probabilistic corrections to account for mass accreted by recycled pulsars in binary systems to mass measurements of 90 neutron stars, we find that the birth masses of neutron stars can be described by a unimodal distribution that smoothly turns on at $\mathbf{1.1\,\mathrm{\mathrm{M}_{\odot}}}$, peaks at $\mathbf{\sim 1.27\,\mathrm{\mathrm{M}_{\odot}}}$, before declining as a steep power law. Such a “turn-on” power-law distribution is strongly favoured against the widely adopted empirical double-Gaussian model at the $\mathbf{3\sigma}$ level. The power-law shape may be inherited from the initial mass function of massive stars[1, 2], but the relative dearth of massive neutron stars implies that single stars with initial masses greater than $\mathbf{\sim 18\,\mathrm{\mathrm{M}_{\odot}}}$ do not form neutron stars, in agreement with the absence of massive red supergiant progenitors of supernovae[3, 4].
@@
Henan Academy of Sciences, Zhengzhou 450046, Henan, China
@@
Department of Physics and Synergetic Innovation Center for Quantum Effects and Applications, Hunan Normal University, Changsha, China.
@@
Institute of Interdisciplinary Studies, Hunan Normal University, Changsha, China
@@
Centre for Gravitational Astrophysics, College of Science, The Australian National University, Canberra, Australian Capital Territory, Australia.
@@
Institute for Globally Distributed Open Research and Education (IGDORE), Gothenburg, Sweden.
@@
Australia Telescope National Facility, CSIRO Space and Astronomy, Epping, New South Wales, Australia.
School of Physics and Technology, Wuhan University, Wuhan, Hubei 430072, China
@@
The observational determination of mass function of neutron stars is a long-standing problem in astrophysics. Early attempts were able to derive only loose constraints on the range of possible neutron star masses from a small number of measurements[5, 6]. For a long time, all observed neutron-stars masses were in a narrow range, consistent with a Gaussian distribution with a mean of $1.35\,\mathrm{\mathrm{M}_{\odot}}$ and a width of $0.04\,\mathrm{\mathrm{M}_{\odot}}$ (ref.[7]). Owning partly to the precise mass measurements of the Hulse-Taylor binary pulsar[8, 9], a canonical mass of $1.4\,\mathrm{\mathrm{M}_{\odot}}$ has routinely been adopted in most current studies of neutron stars and in textbooks.
@@
As the number of neutron-star mass measurements steadily increased, especially with the discoveries of massive neutron stars around $2\,\mathrm{\mathrm{M}_{\odot}}$ (refs.[10, 11]), it became apparent that the single-Gaussian model could no longer fit all available data. Various groups have started to employ a two-Gaussian (2G) model to explain the observations, with the secondary mass peak shifting to higher values as more massive neutron stars have been measured[12, 13, 14, 15]. The most widely adopted model consists of a narrow peak at $1.4\,\mathrm{\mathrm{M}_{\odot}}$ and a broad peak at $1.8\,\mathrm{\mathrm{M}_{\odot}}$ (ref.[16]), and the two peaks are attributed to different formation channels and evolutionary histories of the neutron stars[17].
@@
From a theoretical perspective, the neutron-star mass function is a powerful probe into supernova physics, compact binary evolution and the neutron-star equation of state. It was once proposed that the two low-mass peaks (at $1.25\,\mathrm{\mathrm{M}_{\odot}}$ and $1.35\,\mathrm{\mathrm{M}_{\odot}}$) claimed in the mass distribution are the product of distinct supernova progenitor channels, electron-capture supernovae and low-mass iron-core-collapse supernovae[18, 19]. An additional high-mass peak centred around $1.8\,\mathrm{\mathrm{M}_{\odot}}$ in the mass distribution was thought to arise due to the transition from convective central carbon burning to radiative central carbon burning[18, 20]. Recent detailed studies, however, have revealed a complicated landscape of interacting shells with discontinuous changes in the core and remnant masses when the number of shells changes, or shells merge, as a function of initial or helium core mass[21, 23].
@@
Efforts to link the observed mass distribution to neutron star formation channels are complicated because the fact that neutron stars in binary systems often go through a recycling process and gain some mass through accretion[24, 25, 26, 27]. Such a recycling process is the standard formation pathway for millisecond pulsars, which have by far dominated the population of neutron stars with precise mass measurements. Although the recycling process may increase the neutron-star mass by up to $\sim 0.3$ M (refs.[24, 27, 28]), it has been argued that the high-mass peak around $1.8\,\mathrm{\mathrm{M}_{\odot}}$ in the observed mass spectrum is probably not caused by mass accretion but, instead, a hint of neutron stars being born massive[15]. Current supernova simulations, however, have difficulties in producing enough neutron stars around $1.8\,\mathrm{\mathrm{M}_{\odot}}$ as implied by the empirical 2G model[29].
@@
To determine the neutron-star mass distribution, we compiled a sample of 90 neutron stars for which well-determined mass estimates are available from observations of radio pulsars, gravitational waves and X-ray binaries. For most of the neutron stars in our sample, the observed spin properties allowed us to securely categorize them into two subclasses: recycled and non-recycled (slowly rotating) neutron stars. The first class is characterized by rapid spins (period $P\lesssim 100\,\mathrm{ms}$), whereas the second class includes neutron stars with spin periods of the order of seconds or longer. When constraining the birth-mass function of neutron stars, the key difference between the two subclasses is that observed masses of recycled pulsars need to be corrected for mass accreted throughout the recycling process, whereas the measured masses of slow neutron stars should equal their birth masses, as no mass-gaining process occurred. See Section A, Extended Data Tables 1 and 2 for details.
@@
Figure 1 shows the distributions of measured masses (black lines with grey-shaded areas) and estimated birth masses (blue lines) for 39 recycled pulsars. In each subplot, we also list the pulsar name, observed spin period and plausible initial spin period. The birth masses of recycled pulsars were estimated by subtracting accreted masses from their measured masses. We adopted an analytical approach for estimating the accreted mass. To the leading order, the total amount of accreted mass scales inversely with the final equilibrium spin period in the recycling process. Based on the simple accretion spin-up theory[24, 30, 27], we derived the probability distribution of the accreted mass by accounting for uncertainties in the accretion rate, neutron star moment of inertia, magnetic inclination angle and accretion disk-magnetosphere interactions. See Section B and Extended Data Figs 1,  2 and  3.
@@
We fitted a variety of models to the mass distribution of neutron stars. The Gaussian models are a 2G distribution with both a low- and high-mass cutoff ($\rm 2G^{max}_{min}$ ), a 2G distribution with a low-mass cutoff ($\rm 2G_{min}$) and a 2G distribution with a high-mass cutoff ($\rm 2G^{max}$). In addition to the Gaussian-family models, we considered a turn-on power-law(TOP) distribution, which allows an exponential turn-on from the minimum mass[31], as well as a skewed Student’s $t$-distribution. Some of the models are illustrated in panel b of Figure 2. Bayesian model selection was performed to find the model that best describes the data. The Bayes factors, which quantify the statistical support for one model against another, between five top-performing models and the empirical two-Gaussian (2G) model are shown in panel a of Figure 2. The best-performing model was the turn-on power law (TOP), which is preferred against the 2G model with a Bayes factor greater than 300. Assuming equal prior probabilities for both models, this result means an odds of $0.003$, or equivalently $\sim 3\sigma$ significance. Potential selection effects are discussed in Section G of Methods, and further details can be found in Extended Data Figure 6.
@@
Figure 3 shows the reconstructed birth-mass function of neutron star for the turn-on power law model, as well as histograms for two different bin sizes. We found that the minimum mass was $1.1^{+0.04}_{-0.05}\,\mathrm{\mathrm{M}_{\odot}}$. The distribution peaked at $1.27^{+0.03}_{-0.04}\,\mathrm{\mathrm{M}_{\odot}}$ and declined with a power-law index of $6.5^{+1.3}_{-1.2}$. We found marginal evidence for a maximum-mass cutoff, with a Bayes factor of 2. The maximum mass was loosely constrained to be $2.36^{+0.29}_{-0.17}\,\mathrm{\mathrm{M}_{\odot}}$. (All numbers quoted are median and $1\sigma$ credible intervals.) The posterior distributions of model parameters are presented in Figure 4. We found no evidence for other peaks on top of the power-law function, although such a possibility cannot be ruled out either.
@@
The final evolution of the lowest-mass stars that make iron-core-collapse supernovae is notably diverse and a challenge to model due to the complicated physics[32]. Hence, any constraints from observations on this mass regime are invaluable. Such low-mass iron-core-collapse supernovae are predicted to produce neutron stars with masses as low as or potentially below $1.15\,\mathrm{M}_{\odot}$ (refs.[18, 32]), consistent with our observed minimum. Electron-capture supernovae are unlikely to produce such low-mass neutron stars[33]. They, however, may make neutron stars with masses within the peak of the distribution. The data presently do not provide support for the two discrete low-mass peaks that are expected from theoretical calculations[19]. Instead, the distribution is probably a superposition of several overlapping discontinuous (as a function of initial mass) contributions[21, 23], overlaid with some scatter due to variations in the turbulent supernova explosions[34]. Particularly marked is the absence of the second peak predicted at $1.35\,\mathrm{\mathrm{M}_{\odot}}$, which probably hints at the larger importance of lower-mass core-collapse and of stripped-core supernova progenitors[35].
@@
If we accept the core structures predicted by current stellar evolution models, the absence of prominent peaks and the measured power-law slope of the neutron-star mass distribution have important implications for supernova explosion physics. Specifically, the most parsimonious explanation of TOP is that neutron stars from more massive progenitors undergo substantial net accretion during the explosion and that explosions occur only for stars with main sequence masses of $\lesssim 18\,\mathrm{\mathrm{M}_{\odot}}$ in agreement with the observed absence of massive red supergiants as progenitors for core-collapse supernovae[3, 4] or for stripped stars with similar core structures. We demonstrate this through the use of semi-analytic supernova models[21], as described in Methods Section F.
[…]
@@
Figure 1: Individual mass distributions of 39 recycled pulsars. Black lines with grey shaded area are measured masses, whereas blue lines are estimated birth masses after subtracting accreted masses based on the analytical prescription. Observed and plausible initial spin periods, along with pulsar names are given in each subplot. For readability, we scale different posteriors to the same ${y}$ coordinate height.
@@
Figure 2: Example mass models and their Bayes factors. a, Bayes factor between various models (M) and the 2G model for the birth mass function of neutron stars. b, Mass models. In addition to the models listed in a, we also show a single-Gaussian distribution with both a low- and high-mass cutoff ($\rm G^{max}_{min}$). Model parameters were set to their maximum-likelihood values inferred from data. SST, skewed Student’s $t$-distribution.
@@
Figure 3: Reconstructed birth-mass function of neutron stars based on the preferred TOP model. The solid blue line shows the population predictive distribution, whereas the shaded blue band indicates the 90% credible interval. Also plotted are a ten-bin histogram (mean value) with 90% credibility error bars and a 20-bin histogram (mean value) for the data, showing the robustness of our model in fitting the underlying distribution. CI, confidence interval.
@@
Figure 4: Posterior distribution of the TOP model parameters. Shown are the minimum neutron-star mass $m_{\rm min}$, the peak of the mass function $m_{\rm peak}$, the maximum mass $m_{\rm max}$ and the power-law index $\alpha$.
@@
All the neutron-star mass measurements used in this study are listed in Extended Data Tables 1 and 2 with the original references. These mass measurements, accreted-mass corrections for recycled pulsars, posterior samples from Bayesian inference and the data behind Extended Data Tables 1 and 2 and Figures 1–4 and Extended Data Figures 1–4 are available via Zenodo at https://doi.org/10.5281/zenodo.14375273. Source data are provided with this paper.
The following open-source software packages were used in this paper: GalDynPsr, gwpopulation, BILBY and dynesty. The Python scripts used for data analysis and figure generation are publicly available via GitHub at https://github.com/GW-BNUZ/NSbirthMass.
We thank Z. Li, X. Chen and Z. Han for providing data from their numerical simulations, which are plotted in Extended Data Figure 2. We also thank I. Mandel and M. Bailes for useful discussions. This work was supported by the National Natural Science Foundation of China (Grant Nos. 12203004 to X.Z., 12021003 and 12433001 to Z.-H.Z., 12305059 to Z.-Q.Y. and 12405056 to Z.-C.C.). X.Z. and H.G. are supported by the Fundamental Research Funds for the Central Universities. This work was supported in part by the Australian Research Council (ARC) Centre of Excellence for Gravitational Wave Discovery (Project Nos. CE170100004 and CE230100016). S.S. is a recipient of an ARC Discovery Early Career Research Award (DE220100241). This work was supported by ARC Discovery Grants DP240101786 (to B.M. and A.H.) and DP240103174 (to A.H.).
@@
The authors declare no competing interests.